# ETF: An Entity Tracing Framework for Hallucination Detection in Code Summaries

### v3

@@ 5 Methodology
This behaviour aligns with working memory theory in cognitive science Baddeley and Hitch (1994); working memory is a brain system that temporarily stores and manipulates the information necessary for complex cognitive tasks like learning and reasoning. The capacity of working memory is bounded by $7\pm 2$ object at any point in time, which further reduces to 2-3 objects if the objects have relational dependencies with each other. Since code summaries often involve interdependent objects, developers must focus on local aspects to build a global understanding, suggesting a bottom-up heuristic for code summary comprehension.
We leverage these behavioural insights to design an LLM-powered framework for detecting hallucinations in code summaries, which involves tracing the entities from the summary to the code. This aspect of mapping the entities from the summary to the code aims to simulate the bottom-up behavioural model of verifying the description of coding entities at a time. With these insights, we aim to measure the correctness of a code summary as a two-step process: (1) Entity Verification and (2) Entity-Intent Verification. The detailed flow of this framework can be found in Figure 1.
@@ 5.1 Entity Verification
In entity verification, we check if the entities in the summary are present in the source code to detect extrinsic hallucination. This involves extracting entities from both the code and summary, then mapping entities from summary to the code. We elaborate on this process below:
Entity Extraction from code: We leverage program analysis to extract entities from code (Javalang Python package  [^3] https://github.com/c2nes/javalang). The code is tokenized (lexer) and parsed into an abstract syntax tree (AST). This tree structure represents the hierarchical organization of code elements, making it easier to analyze. This yields a fine-grained classification of all the tokens present in the code such as variable names, class names, function names, etc.
Entity Extraction from summary: Tabassum et al. (2020) propose the task of entity detection in code summaries and introduce a relevant NER tagset. We adopt this tagset for extracting entities from code summaries (Prompt: Appendix A Figure 5).
Leveraging LLMs to recognize entities introduces the risk of hallucinations, where the model may fabricate entities not present in the code summary. To address this, we implement a filtration step to remove such fabricated entities. We evaluate Gemini and GPT-4-Omni for Code NER using human-collected data, with results in Appendix C. Additionally, we assess an open-source model as part of our contribution. Our findings show a strong correlation between GPT-4-Omni predictions and human data, confirming its effectiveness for entity detection in our framework.
@@ 5.2 Entity-Intent Verification
The presence of an entity in both the summary and code indicates that the entity is valid but does not warrant the correctness of the context in which it is discussed. For example, in Figure 1 jobId is a correct entity, but the context of retrieving jobID from the database is incorrect. To address this problem, we propose verifying whether the intent of each mapped entity is accurately described in the summary. We extract all sentences containing the entity of interest from the summary to form its intent context. To identify these relevant sentences that describe an entity’s intent, we explored two approaches: (1) prompting a language model to find relevant sentences and (2) using string-matching heuristics (Refer Appendix D). Qualitative assessment showed that rule-based heuristics were more effective and efficient than prompt-based methods, which suffered from hallucinations. Therefore, we relied on string-matching-heuristics for our framework. After identifying the entity and intent, we use LLMs with zero-shot prompting to verify their correctness with the code (Prompt: Appendix 5).We also experimented with few-shot prompting by including examples of various hallucination types in code summaries along with the representative code. However, performance degraded due to the increased prompt length, consistent with findings in recent works like Mirzadeh et al. (2024).
5.3 Instance Level Hallucination Detection
To identify the quality of the whole summary with respect to the code, we aggregate the individual entity-intent hallucination and set the threshold for labelling as 1, as discussed in (Section 1). Specifically, the summary is marked as hallucinated if at least one intent is identified as INCORRECT.
@@ 6 Experiments and Results
For our experiments, we consider summaries from instruction-tuned versions of the SOTA code and language models, from IBM-Granite family (20B-instruct; 34B-instruct) Mishra et al. (2024), Llama3 family (8B-instruct and 70B-instruct) Touvron et al. (2023), CodeLlama family (7B and 34B) Roziere et al. (2023) and Mistral family (7B-instruct) Jiang et al. (2023).
For intent verification, we started our initial experiments with powerful open-source models like LLama3-70B Touvron et al. (2023) but observed worse than random performance (F1 score- 0.37) due to the very high complexity of the task. Similar qualitative studies with other models in our preliminary experiments prompted us to explore advanced models with stronger capabilities in code reasoning, such as GPT4-Omni Achiam et al. (2023) and Gemini-1.5-Flash Team et al. (2023). Hyperparameters and other details can be found in Appendix (B).
@@
Gemini-1.5-Direct 0.41 0.49 0.25
Gemini-1.5-ETF* 0.68 0.64 0.51
@@
Table 2: Hallucination Detection results. We consider two evaluation aspects: 1) Instance Level, which aims to label the entire summary, and 2) Entity level, which labels individual entities in summaries. Here, P, R and F1 refer to the precision, recall and F1 score.
@@
Models Summary Length CE count Fabricated ()($\downarrow$)( ↓ ) Mapped()($\uparrow$)( ↑ ) Unmapped ()($\downarrow$)( ↓ )
Codellama-7B 236.10 8.638 35.68% 80.17% 17.65%
Mistral-7B 227.91 6.961 8.59% 79.92% 17.22%
Llama3-8B 257.21 9.45 6.44% 84.50% 12.77%
Granite-20B 148.95 7.10 4.25% 79.54% 19.64%
Granite-34B 214.50 6.55 2.45% 85.69% 12.41%
Codellama-34B 278.67 8.22 4.87% 79.78% 19.05%
Llama3-70B 313.58 10.01 0.15% 88.44% 8.69%
@@
Table 3: Quantitative Analysis: This table presents summary statistics for all seven models, where ()($\uparrow$)( ↑ ) means a higher score is better and ()($\downarrow$)( ↓ ) means lower is better. "CE count" is the number of code entities in the summary. "Fabricated" is the percentage of fabricated entities during named entity recognition, normalized by the total entities. "Mapped" is the percentage of entities correctly mapped to the code, and "Unmapped" refers to those unmapped, both normalized by the total entity count.
@@
Entity-Intent Verification: In this aspect of evaluation, we aim to verify the intent of an individual entity. We report the results of entity-intent verification in the Table 2. It can be observed that the GPT4-Omni F1-Score is 0.61 while the Gemini F1 Score is 0.55. Upon analysis, we found that these models often classify INCORRECT tuples as CORRECT. This is mainly due to the convincing nature of the summary, which may also be subjective due to a lack of proper code context. For instance, we observe significant errors when the code references a function or library that is not defined in the input. In such cases, the model infers the functionality based on the library name (Identifier Name Bias 4), which can be difficult to verify.
@@
Instance Level Hallucination Verification: In this aspect of evaluation, we aim to verify the overall summary instance. To compare our approach, we consider a direct setup which involves providing a <code, summary> tuple to identify if the summary is hallucinated or not. We provide these results in Table 2, and it can be observed that our approach provides significant improvement in F1-Score when compared to the Direct approach. In general, the direct evaluation method suffers from hallucinations, such as when identified entities for hallucination are absent from the summary or when natural language entities are mistakenly considered code entities, overall resulting in poor performance. This conveys that our finer-grained evaluation provides a reliable method to identify hallucinated summaries. It also helps with interpretability as it identifies the hallucinated sections of the summary.
@@ 7 Analysis
In this section, we discuss various quantitative and qualitative insights of our framework. We first discuss summaries generated by individual models and then elaborate on the general predictive behaviour of our framework and cases of errors.
@@ 7.2 Predictive Analysis
Our framework captures ungrounded entities in summaries. For example, when the model refers to a non-existent Hub API based on the keyword ‘HubException", the unmapped entities such as "Hub" are identified. This enables fine-grained and interpretable detection of hallucinations. We can see that the ungrounded entities like "Hub" are captured here, leading to more fine-grained and interpretable detection of hallucination.
[ht!] private List<Transaction> retrieveTransactions(String rowStatusCd) throws HubException {................} Summary: This method retrieves transactions using the “Hub" API based on the input parameters by the “Hub database". It interacts with several key entities, including “squid:s1166" and “squid:s1172", to gather the data…… Analysis: Hallucination and Unmapped Entities
@@ 7.3 Error Analysis
This section discusses the two major error cases in our framework:
Error Case 1-Creative Summary:
public static int writeShort(ArrayView target, int offset, short value) {...} Summary:  Here’s a more detailed implementation of the above method: ….{JAVA CODE}….. Creative Summary by Mistral 7B
During the generation, the model may mention certain aspects of the code in a creative way which may not be incorrect. In the given example, the code summary discusses a more elaborate version of the input code by restating an elongated version. Here, the entities present in the summary are predicted to be ungrounded by our framework.
@@
protected PreparedStatement setObjects(Connection connection, String query, Object... objects) throws SQLException {....} Summary:  ……set the parameters of the prepared statement using the provided objects. This method is likely responsible for iterating over the objects and setting them as parameters in the prepared statement. Finally, the method returns the prepared statement object….. Changed Entity Form by Llama3-70B
@@ 8 Conclusion and Future Work
Our work addresses the critical challenge of detecting hallucinations in code summarization, a task that demands a deep understanding of both programming and natural languages. By introducing a novel dataset and the Entity Tracing Framework (ETF) with 73% F1 score, we establish a systematic approach to grounding code entities within summaries, enabling a more interpretable and accurate evaluation of explanations. In the future, this framework can be enhanced by incorporating a multi-agent system and leveraging multiple LLMs in tandem to improve prediction accuracy. The current framework can be further developed to better mitigate the occurrence of hallucinations.
@@ 9 Limitations
While the framework is designed to be generic, certain components—such as code parsers for entity detection—may be unavailable for low-resource programming languages like COBOL or Perl. Moreover, the performance of large language models (LLMs) heavily depends on their parameter size and the volume of training data. Given the complexity of this task, smaller open-source models may struggle to perform effectively, reinforcing the need for larger LLMs. However, these larger models may not be scalable and often demand significantly greater computational resources.
@@ Appendix A Prompts
Intent Verification Prompt Assume you are an expert in understanding JAVA code. Your task is to verify whether the description of ’mapped_entity’ in the given text is correct, incorrect, or irrelevant with respect to the code. Only output one of the following labels: [“CORRECT", “INCORRECT", “IRRELEVANT"]. Description: $\{relevant\_sent\}$ $[CODE]$CODE} $[\mathbin{/}CODE]$
@@
Named Entity Recognition Prompt Assume you are an expert in understanding Java and performing named entity recognition related to Java code. You have to label the entities by considering the following labels: Code Entities: CLASS, VARIABLE, FUNCTION, LIBRARY, VALUE, DATA TYPE, and HTML or XML TAG Natural Language Entities: APPLICATION, UI ELEMENT, LANGUAGE, DATA STRUCTURE, ALGORITHM, FILE TYPE, FILE NAME, VERSION, DEVICE, OS, WEBSITE, and USER NAME. For every entity in the input mention the entity_type in the given format only. Strictly follow this template and only print the output without any other word. You can follow the example below: “‘ {Incontext Example} “‘ Now consider the summary describing a code below: {generated_summary}
@@
Direct Evaluation Prompt Assume you are an expert in understanding JAVA code. Your task is to verify if the description of the code entities present in the given summary is correctly described or NOT as per the code logic. Output all the ‘entity_name’ and a relevant_sentence’ corresponding to the ‘entity_name’, which are incorrectly described. Do not provide any other details. Strictly follow this format: $[{entity\_name:``",relevant\_sentence:``"}]$ Summary: {SUMMARY} Code: {CODE}
@@ Appendix C NER Evaluation
This section discusses the NER performance of various models considered in this work. To perform NER using LLMs, we provide the code summary and NER tagset in the prompt (Appendix 5) using a one-shot in-context example to extract all the entities discussed in the summary accompanied by their types. To evaluate the entity extraction, we assess two key aspects: entity coverage and entity type correctness. 1) Entity Coverage: This measures whether all valid entities in the summary are detected. We quantify this using the Jaccard Similarity between the entities in the generated output and those in the ground truth. 2) Entity Type Correctness: This evaluates whether the detected entities have been assigned the correct types. For this, we use the F1 score as the metric.
@@
Models Jaccard Similarity F1
Llama3-8B 0.5298 0.78
Llama3-70B 0.5981 0.90
Mistral-7B 0.4458 0.65
Granite-20B 0.4897 0.85
Granite-34B 0.48181 0.84
Codellama-7B 0.4586 0.84
Codellama-34B 0.5079 0.83
Table 5: NER Results on GPT Data
Appendix D String Matching Heuristics

### v4

@@ 5 Methodology
This behaviour aligns with working memory theory in cognitive science Baddeley and Hitch (1994); working memory is a brain system that temporarily stores and manipulates the information necessary for complex cognitive tasks like learning and reasoning. The capacity of working memory is bounded by $7\pm 2$ objects at any point in time, which further reduces to 3-4 objects Cowan (2014) if the objects have relational dependencies with each other. Since code summaries often involve interdependent objects, developers must focus on local aspects to build a global understanding, suggesting a bottom-up heuristic for code summary comprehension. We leverage these behavioural insights to design an LLM-powered framework for detecting hallucinations in code summaries, which involves tracing the entities from the summary to the code. This aspect of mapping the entities from the summary to the code aims to simulate the bottom-up behavioural model of verifying the description of coding entities at a time. With these insights, we aim to measure the correctness of a code summary as a two-step process: (1) Entity Verification and (2) Entity-Intent Verification. The detailed flow of this framework can be found in Figure 1.
@@ 5.1 Entity Verification
In entity verification, we check if the entities in the summary are present in the source code to detect extrinsic hallucination. This involves extracting entities from both the code and summary, then mapping entities from the summary to the code. We elaborate on this process below:
Entity Extraction from code: We leverage program analysis to extract entities from code (Javalang Python package555javalang Repository). The code is tokenised (lexer) and parsed into an abstract syntax tree (AST). This tree structure represents the hierarchical organisation of code elements, making it easier to analyse. This yields a fine-grained classification of all the tokens present in the code, such as variable names, class names, function names, etc.
Entity Extraction from summary: Tabassum et al. (2020) proposes the task of entity detection in code summaries and introduces a relevant NER tagset. We adopt this tagset for extracting entities from code summaries (Prompt: Appendix A Figure 5). Leveraging LLMs to recognise entities introduces the risk of hallucinations, where the model may fabricate entities not present in the code summary. To address this, we implement a filtration step to remove such fabricated entities. We evaluate Gemini and GPT-4-Omni for Code NER using human-collected data, with results in Appendix C. We also assess all the open-source models considered in this study. Our findings show a strong correlation between GPT-4-Omni predictions and human data, confirming its effectiveness for entity detection in our framework.
@@ 5.2 Entity-Intent Verification
The presence of an entity in both the summary and code indicates that the entity is valid, but does not warrant the correctness of the context in which it is discussed. For example, in Figure 1 jobId is a correct entity, but the context of retrieving jobID from the database is incorrect. To address this problem, we propose verifying whether the intent of each mapped entity is accurately described in the summary. We extract all sentences containing the entity of interest from the summary to form its intent context. To identify these relevant sentences that describe an entity’s intent, we explored two approaches: (1) prompt-based and (2) string-matching heuristics. Our qualitative assessment, detailed in Appendix D, demonstrates that rule-based heuristics were both more effective and efficient than prompt-based methods, which were prone to hallucinations. Therefore, we relied on string-matching heuristics for our framework. After identifying the entity and intent, we use LLMs with zero-shot prompting to verify their correctness with the code (Prompt: Appendix 5). We also experimented with few-shot prompts by including examples of various hallucination types in code summaries along with the representative code. However, performance degraded due to the increased prompt length, consistent with findings in recent works like Mirzadeh et al. (2024).
To identify the quality of the whole summary with respect to the code, we aggregate the individual entity-intent hallucination and set the threshold for labelling as 1, as discussed in (Section 1) following human annotation where a summary was rated as ’FAIR’ or ’POOR’ when an average of 1.33 entities were wrongly described.
@@ 6 Experiments and Results
Gemini-2.0-Direct 0.51 0.50 0.42
Gemini-2.0-ETF* 0.64 0.65 0.64
@@
Mixtral-8x22B-Direct 0.48 0.48 0.45
Mixtral-8x22B-ETF* 0.62 0.61 0.61
Llama-3.1-70B-Direct 0.57 0.54 0.38
Llama-3.1-70B-ETF* 0.62 0.62 0.54
Llama3-8B-Direct 0.59 0.51 0.26
Llama3-8B-ETF* 0.51 0.55 0.50
Mistral-7Bv3-Direct 0.16 0.50 0.24
Mistral-7Bv3-ETF* 0.51 0.50 0.41
@@
Table 2: We report macro Precision (P), Recall (R) and F1. We consider two evaluation aspects: 1) Instance Level, which aims to label the entire summary, and 2) Entity Level, which labels individual entities in summaries. Here, * represents the proposed approach.
For summary generation, we consider instruction-tuned versions of the IBM-Granite family (20B and 34B) Mishra et al. (2024), Llama3 family (8B and 70B) Touvron et al. (2023), CodeLlama family (7B and 34B) Roziere et al. (2023) and Mistral-7B Jiang et al. (2023).
For intent verification, we consider Llama3.1-70B, Llama3-8B Touvron et al. (2023), Mixtral-8x22B Jiang et al. (2024a), Mistral7B-v3 Jiang et al. (2023), GPT4-Omni Achiam et al. (2023) and Gemini2-Flash Team et al. (2023). All experimental details can be found in Appendix B.
@@
Models Summary Length CE count Mapped $(\uparrow)$ Unmapped $(\downarrow)$ NL entities
Codellama-7B 236.10 8.638 80.17% 17.65% 2.18%
Mistral-7B 227.91 6.961 79.92% 17.22% 2.86%
Llama3-8B 257.21 9.45 84.50% 12.77% 2.73%
Granite-20B 148.95 7.10 79.54% 19.64% 0.82%
Granite-34B 214.50 6.55 85.69% 12.41% 1.90%
Codellama-34B 278.67 8.22 79.78% 19.05% 1.17%
Llama3-70B 313.58 10.01 88.44% 8.69% 2.87%
@@
Table 3: Quantitative Analysis: This table presents summary statistics for all seven models. "CE count" is the number of code entities in the summary. "Mapped" is the percentage of code entities correctly mapped to the code, "Unmapped" refers to those unmapped, and "NL entities" refers to the remaining natural language (non-code) entities, all normalised by the total entity count.
@@
Entity-Intent Verification: In this aspect of evaluation, we aim to verify the intent of an individual entity. We report the results of entity-intent verification in the Table 2. In general, we observe consistent improvements across all the models compared to the direct approach. It can be observed that the GPT4-Omni F1-Score is 0.73 while the Gemini F1 Score is 0.64. We observed close to random performance for other models like Llama3.1-70B, Llama3-8B and Mistral-7 B. Upon analysis, we found that these models often classify INCORRECT tuples as CORRECT when the code references a function or library that is not defined in the input. Here, the model infers the functionality based on the library name (Identifier Name Bias 4), which can be difficult to verify.
@@
Instance Level Hallucination Verification: In this aspect of evaluation, we aim to verify the overall summary instance. To compare our approach, we consider a direct setup which involves providing a <code, summary> tuple to identify if the summary is hallucinated or not (Refer Appendix A, Figure 7). We provide these results in Table 2, and it can be observed that our approach provides significant improvement in F1-Score when compared to the Direct approach. In general, the direct evaluation method suffers from hallucinations, such as when identified entities for hallucination are absent from the summary or when natural language entities are mistakenly considered code entities, overall resulting in poor performance. This conveys that our finer-grained evaluation approach, which traces the entities from summary to code, provides a more reliable method to identify hallucinated summaries compared to traditional LLM-based approaches.
@@ 7 Analysis
In this section, we discuss various quantitative and qualitative insights of our framework, including its general predictive behaviour and cases of errors.
@@ 7.1 Quantitative Analysis
7.2 Taxonomy Analysis
In this section, we discuss insights related to each category defined in the proposed taxonomy. As described in the Figure 3, we observe Code Complexity to be the most common hallucination cause. This arises due to the complicated variable names, which are not suggestive of what the code is about, the lengthy code or the complex logic. This may also stem from a lack of reasoning capability in models, often due to their small size or insufficient training. Some of the most confused cases in our dataset involve function overloading and recursive functions, where the models often got confused between the caller function and the callee function. Other scenarios include: (1) the presence of multiple conditions within a boolean expression (e.g., an if condition with a long boolean argument containing multiple operators and operands), (2) multiple function calls within a single expression, and (3) the use of variable names that lack meaningful semantic interpretation.
The next major category is Insufficient Knowledge, where models fail to correctly interpret fundamental aspects of the code or the libraries used. This includes cases where models cannot recognise function arguments or fail to understand library functionalities, often due to limited exposure during training or user-specific customisations of the libraries. Such hallucinations are more prevalent in smaller models, though occasional instances are also observed in larger models (greater than 30B).
In addition to the above categories, we also observe notable instances of Name Bias and hallucination influenced by Natural Language Context. Primarily, Name Bias occurs when the models rely too much on the semantic interpretation of the identifier names. This often overlaps with Insufficient Knowledge, where the model lacks grounding but guesses based on naming, and can be attributed to inadequate model training. For example, there might be a line which mentions a function name not defined in the given code itself; in such scenarios, models are ‘forced’ to rely on the semantics. This is also the reason that there are some frequent cases of Insufficient knowledge and Name Bias together. It is noteworthy to mention that almost all the annotators also expressed ambiguity in such instances, making it difficult to annotate even for humans. Hallucination due to natural language context, like comments or log statements, occurs when the comments represent outdated information or when comments discuss some additional functionality not suggested in the code logic.
@@ 7.3 Predictive Analysis
[ht!] private List<Transaction> retrieveTransactions(String rowStatusCd) throws HubException{............} Summary: This method retrieves transactions using the “Hub" API based on the input parameters by the “Hub database". It interacts with several key entities, including “squid:s1166" and “squid:s1172", to gather the data…… Analysis: Hallucination and Unmapped Entities
Our framework captures ungrounded entities in summaries. For example, when the model refers to a non-existent Hub API based on the keyword ‘HubException", the unmapped entities such as "Hub" are identified. This enables fine-grained and interpretable detection of hallucinations.
@@ 7.4 Error Analysis
This section discusses the two major error cases which can be observed in our framework:
Error Case 1-Creative Summary: Models sometimes add elaborative details that may be accurate but not explicitly present in the code. In the given example, the code summary discusses a more elaborate version of the input code by restating an elongated version. Here, the entities present in the summary are predicted to be ungrounded by our framework.
public static int writeShort(ArrayView target, int offset, short value) { return writeShort(target.array(), target.arrayOffset()+offset, value); } Summary:  Here’s a more detailed look at the implementation: ….{JAVA CODE}….. Creative Summary by Mistral 7B
@@
protected PreparedStatement setObjects(Connection connection, String query, Object... objects) throws SQLException { PreparedStatement ps = statement(connection, query); setObjects(ps, objects); return ps; } Summary:  ……The method is used to create and execute a prepared statement using the given connection and query. The method takes an array of objects as input, which are used to set the parameters in the prepared statement. The method returns the prepared statement object that was created and executed…… Changed Entity Form by Codellama 7B
@@ 8 Conclusion and Future Work
In this work, we address the problem of detecting hallucinations in code summarisation, a task that demands a deep understanding of both programming and natural languages. By introducing a novel dataset, CodeSumEval and an Entity Tracing Framework (ETF), we present a promising approach to grounding code entities within summaries. This enables a more explainable and accurate evaluation of code summaries. In the future, we plan to explore agent collaboration where different LLMs verify entities and intents independently. Further improvements may involve fuzzy or embedding-based entity detection, documentation lookup or extension to mitigate hallucinations.
@@ 9 Limitations
While the framework is designed to be generic, certain components, such as code parsers for entity detection, may be unavailable for low-resource programming languages like COBOL or Perl. This limits its applicability to low-resource programming languages (e.g., COBOL, Perl) lacking robust static analysis tools. Further, the entity detection and verification stages depend heavily on large-scale LLMs. Given the complexity of this task, smaller open-source models may struggle to perform effectively, reinforcing the need for larger LLMs. This may not be scalable and often demands significantly greater computational resources. Finally, our multi-stage verification process involves multiple LLM calls, making real-time deployment and integration into production-level developer tools challenging without further optimisation.
@@ Appendix A Prompts
Intent Verification Prompt Assume you are an expert in understanding JAVA code. Your task is to verify whether the description of ’mapped_entity’ in the given text is correct, incorrect, or irrelevant with respect to the code. Only output one of the following labels: [“CORRECT", “INCORRECT", “IRRELEVANT"]. Description: $\{relevant\_sent\}$ $[CODE]$ {CODE} $\mathbin{/}CODE]$
@@
Named Entity Recognition Prompt Assume you are an expert in understanding Java and performing named entity recognition related to Java code. You have to label the entities by considering the following labels: Code Entities: CLASS, VARIABLE, FUNCTION, LIBRARY, VALUE, DATA TYPE, and HTML or XML TAG Natural Language Entities: APPLICATION, UI ELEMENT, LANGUAGE, DATA STRUCTURE, ALGORITHM, FILE TYPE, FILE NAME, VERSION, DEVICE, OS, WEBSITE, and USER NAME. For every entity in the input, mention the entity_type in the given format only. Strictly follow this template and only print the output without any other words. You can follow the example below: “‘ {Incontext Example} “‘ Now consider the summary describing the code below: {generated_summary}
@@
Direct Evaluation Prompt Assume you are an expert in understanding JAVA code. Your task is to verify if the description of the code entities present in the given summary is correctly described or NOT as per the code logic. Output all the ‘entity_name’ and a relevant_sentence’ corresponding to the ‘entity_name’, which are incorrectly described. Do not provide any other details. Strictly follow this format: [entity_name:,relevant_sentence:][{entity\_name:``",relevant\_sentence:``"}] Summary: {SUMMARY} Code: {CODE}
@@ Appendix C NER Evaluation
This section discusses the NER performance of various models considered in this work. To perform NER using LLMs, we provide the code summary and NER tagset in the prompt (Appendix 5) using a one-shot in-context example to extract all the entities discussed in the summary, accompanied by their types. To evaluate the entity extraction, we assess two key aspects: entity coverage and entity type correctness. 1) Entity Coverage: This measures whether all valid entities in the summary are detected. We quantify this using the Jaccard Similarity between the entities in the generated output and those in the ground truth. 2) Entity Type Correctness: This evaluates whether the detected entities have been assigned the correct types. For this, we use the F1 score as the metric.
@@
Models JS F1 Fabricated $(\downarrow)$
Codellama-7B 0.4586 0.84 35.68%
Mistral-7B 0.4458 0.65 8.59%
Llama3-8B 0.5298 0.78 6.44%
Granite-20B 0.4897 0.85 4.25%
Granite-34B 0.48181 0.84 2.45%
Codellama-34B 0.5079 0.83 4.87%
Llama3-70B 0.5981 0.90 0.15%
Table 5: NER Results on GPT Data. Here, ‘JS’ refers to Jaccard Similarity, and ‘Fabricated’ refers to the percentage of fabricated entities identified during named entity recognition, normalised by the total entities as mentioned in Table 3.
Appendix D Intent Detection
In this section, we describe the two distinct approaches for intent detection.
D.1 String Matching Heuristics